OpenStudy (anonymous):

8x -2 (x+1) = 2(3x+1)

OpenStudy (anonymous):

8x-2x-2= 6x=2 -2 -2 6x=6x /6 /6 x=x so x=1

OpenStudy (anonymous):

sorry i meant x=x so No solution

OpenStudy (anonymous):

Simplifying 8x + -2(x + 1) = 2(3x + 1) Reorder the terms: 8x + -2(1 + x) = 2(3x + 1) 8x + (1 * -2 + x * -2) = 2(3x + 1) 8x + (-2 + -2x) = 2(3x + 1) Reorder the terms: -2 + 8x + -2x = 2(3x + 1) Combine like terms: 8x + -2x = 6x -2 + 6x = 2(3x + 1) Reorder the terms: -2 + 6x = 2(1 + 3x) -2 + 6x = (1 * 2 + 3x * 2) -2 + 6x = (2 + 6x) Add '-6x' to each side of the equation. -2 + 6x + -6x = 2 + 6x + -6x Combine like terms: 6x + -6x = 0 -2 + 0 = 2 + 6x + -6x -2 = 2 + 6x + -6x Combine like terms: 6x + -6x = 0 -2 = 2 + 0 -2 = 2 Solving -2 = 2 This equation is invalid, the left and right sides are not equal, therefore there is no solution.